Cisplatin, an anticancer drug, has the-molecular formula Pt(NHbase3)base2*Cl(base2). What is the mass (in gram) of one molecule ? (Atomic weights : 195, H = 1.0, N = 14, Cl 35.5)
1
Expert's answer
2020-11-17T10:31:04-0500

Pt(NH3)2Cl2

FMM=195+(14*2)+(1*6)+(35.5*2)

=300g/mol

If 6.022*1023 molecules/mol= 300g/mol

1 molecule= 300*1/6.022*1023

=4.982*10-22g